OpenStudy (anonymous):

The ordered pair (1, -5) is a solution for the system x + 2y = -9 and x - y = 6. True OR False

OpenStudy (zzr0ck3r):

just plug in 1 for x and -5 for y in both equations and see if it makes sense.

OpenStudy (anonymous):

so false? cause 1 + 2 - 5 = -2....

OpenStudy (anonymous):

@zzr0ck3r ???

OpenStudy (anonymous):

hey dude......its true just put x=1 y=(-5) x+2y=-9 1+2(-5) 1+(-10) 1-10 = -9 i.e true budddy nd fr another equation x-y =6 1-(-5) 1+5=6 i.e also true

OpenStudy (anonymous):

hope u understand

OpenStudy (anonymous):

oh i see now, thx